OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x + y = - 7 3x + y = 5 3x - y = -11 x + 3y = 3

OpenStudy (anonymous):

I have no idea what that is. If I knew what to do I wouldn't have asked the question...

sam (.sam.):

Find the slope

OpenStudy (anonymous):

donde esta tu problema, donde te quedaste

sam (.sam.):

\[\Huge Slope=\frac{y-y_1}{x-x_1}\]

sam (.sam.):

slope =-3 Pick any coord,(-3, 2) \[y-y_1=m(x-x1)\] y-2=m(x-(-3)) y-2=-3(x+3) y=-3x-9+2 y=-3x-7

OpenStudy (anonymous):

Um.. So what's the answer, exactly?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

the fisrt one you just have to despejate

OpenStudy (anonymous):

Gracias, Tannerfermat.